OpenStudy (anonymous):

Check my answer? ;;What are the x- and y-intercepts of the line? 2x – 3y = 12

OpenStudy (anonymous):

2x – 3y = 12 Replace ‘y’ with 0 2x – 3*0 = 12 2x – 0 = 12 2x/2 = 12/2 X=6 Then do the same thing to solve for y… 2x – 3y = 12 2*0 – 3y = 12 0 – 3y = 12 -3y = 12 -3y/-3 = 12/-3 Y = -4 (x, y) = (6, -4)

OpenStudy (anonymous):

@iambatman